OpenStudy (anonymous):

Which of the following is equivalent to log 125 25 ?

OpenStudy (anonymous):

log25/log125

hartnn (hartnn):

log 25= log (5^2) = 2 log 5 log 125 = log(5 ^ 3)= 3 log5 can u solve further?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

u understood what i did ? how i got 2log5?

OpenStudy (anonymous):

Yes.

hartnn (hartnn):

so \[\frac{\log 25}{\log 125}=\frac{2 \log 5}{3 \log 5}\] now can u get to final answer?

OpenStudy (anonymous):

2/3

hartnn (hartnn):

yes :)